# Penalized Fair Regression for Multiple Groups in Chronic Kidney Disease

### v1

@@ 4.1 Simulation Implementation
We aimed to explore the performance of various score functions constructed using each of the three different group synthesis mechanisms in Section 3.2 with fairness score weights $\alpha\in\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,0.99\}$. Additionally, we studied different numbers of candidate penalty weight set draws $10,20,\dots,100]$ to empirically characterize changes in accuracy and fairness as the number of candidate sets increases. Lastly, we investigated robustness to functional form misspecification by omitting $X_{1}$ as an input variable. Comparison estimators were a baseline unpenalized regression and the fair reductions constrained regression method from agarwal_reductions_2018 as implemented in the fairlearn python package. The exponentiated gradient algorithm was implemented with allowed fairness constraint violations of $0.1$ and $0.01$, and the grid search algorithm was used with constraint weights of $0.25$, $0.5$, and $0.75$. Otherwise, default hyperparameters were used.
@@ 4.2 Simulation Results
Our novel penalized regression methods were evaluated across several key dimensions: behavior of unfairness penalty weights by fairness score weights, understanding penalty weights across various evaluation measures, effect of changes in multi-group unfairness synthesis function, and the tradeoff between fairness and accuracy. First, we characterized the behavior of unfairness penalty weights under the averaged score functions. The random search yielded different unfairness penalty weights for the three groups across different user-specified fairness score weights (Figure 1). Unfairness penalty weights increased, on average, as the fairness score weight $\alpha$ increased. Unfairness penalty weights also varied across groups and simulation settings.
@@
Next, we examine the impact of the choice of multi-group synthesis function on accuracy and different fairness measures. All three methods for synthesis of unfairness measures across groups in the score function — population-weighted mean, group-weighted mean, and maximum — yielded similar results with overlapping interquartile ranges, especially for accuracy and population-weighted mean fairness (Figure 3). The methods only diverged in maximum unfairness, where the population-weighted mean method underperformed and the maximum method overperformed.
@@
We compare the fairness and accuracy results of our penalized method to the fair reductions constrained regression method (Figure 4). This figure also shows the fairness-accuracy frontier, the set of results for which there is no better performance in both fairness and accuracy (liang_algorithmic_2023). The fairness-accuracy frontier for our penalized method was beyond that of the comparison fair reduction methods; the comparison method did not yield estimators that outperformed our method’s estimators in both mean fairness and accuracy. In simulation settings 1 and 2, the penalized methods were clustered around the fairness-accuracy frontier, while the comparison method had more variance. For simulation setting 3, which featured smaller groups, there was more variance across simulations for both methods, with more replications far from the fairness-accuracy frontier. Results for simulation performance studying variation in the number of candidate penalty weight sets and under estimator misspecification are available in the online appendix. Accuracy and unfairness had less variation as the number of candidate penalty weight sets increased and misspecified estimators generally had fairness-accuracy frontiers shifted downward.
@@
5 CHRONIC KIDNEY DISEASE APPLICATION
5.1 Data
We applied our methods in the American Family Cohort dataset derived from the American Board of Family Medicine PRIME Registry (phillips_prime_2017; stanford_center_for_population_health_sciences_afc_2024). These primary care practice electronic health record data contain billing information (including diagnosis and procedure codes), drug information, and physiological patient measurements from laboratory tests and direct clinical observation. Using input variables measured in 2017-2018, we predicted the occurrence of ESRD in 2019. This outcome was defined through either an ESRD diagnosis code or serum creatinine corresponding to eGFR under 15. eGFR was calculated with the CKD-EPI 2021 equation, which does not adjust for race, to more accurately reflect the true physiological state of patients (inker_new_2021).
Predictors were selected based on prior literature (segal_machine_2020; inoguchi_simplified_2022; su_machine_2022; wang_towards_2022). We had several physiologic predictors. The means of diastolic and systolic blood pressure as well as the maximum and mean values for hemoglobin A1C and billirubin were included as predictors, as was the minimum eGFR value. Missingness was accounted for through the inclusion of indicators of missingness as predictors. Indicators for hypertension emergencies, type 2 diabetes, glomerulonephritis, proteinuria, and pyelonephritis, constructed using diagnosis codes, were included as predictors as well. Diagnosis codes used for these measures are included in the online appendix. Finally, binary documented sex, year of birth, race, and ethnicity were available in the American Family Cohort dataset. We included documented sex and age as predictors. Race and ethnicity were used to define the groups of interest in the construction of fairness penalties and metrics, but we intentionally did not include these variables as predictors. We studied fairness for three race and ethnicity groups: Black, Hispanic non-white, and Asian, with some patients identified as both Black and Hispanic non-white as well as Asian and Hispanic non-white. White patients were treated as the reference group. Two groups who face health disparities in CKD had small sample sizes (American Indian and Alaska Native as well as Native Hawaiian and Pacific Islander) in our dataset and were not included as groups for fairness optimization.
We restricted our sample to patients with diastolic and systolic blood pressure measurements in 2017-2018 and to practices with at least one ESRD patient in 2017-2019 for data quality. To target high-risk susceptible patients, the sample was restricted to patients at least 35 years old with at least two chronic kidney disease diagnosis codes, no documented ESRD prior to 2019, and a maximum systolic blood pressure of at least 140 mmHg, indicating stage 2 hypertension, in 2017-2018 (hsu_elevated_2005; cusick_balancing_2025). Finally, we required recorded values of either race or ethnicity. The resulting sample had 67397 patients, 2804 of whom developed ESRD in 2019. Detailed sample characteristics are described in the online appendix.
5.2 Implementation
We used a decision threshold of 0.15 for estimated ESRD probability, informed by evidence from prior settings that the benefits of increasing the intensity of CKD treatment outweigh the costs and thus that false negatives are more harmful than false positives, but we recognize that other thresholds could also be valid (tisdale_cost-effectiveness_2022). Penalty weight candidates were sampled from the distribution $\text{LogUnif}(-4,-1)$. We used three-fold cross validation to simultaneously identify the penalty weight sets that maximized each score function variation described in Section 4.1. Comparison estimators were the same as in our simulation with the same algorithms and hyperparameters described in Section 4.
5.3 Results
@@
Under the 30 different score function formulations, 6 distinct sets of penalty weights that maximized the score functions and their corresponding penalized regression estimators were identified (Table 2). In most cases, more than one combination of synthesis method and fairness score weight yielded identical results. Overall, the penalized estimators had 0-4 percentage point decreases in accuracy with 1-7 percentage point decreases in unfairness. The penalized regression estimator with group-weighting and fairness score weight 0.1 had the same performance as the population-weighted estimator with fairness score weight 0.1 or 0.2. These three estimators had less than a percentage point difference in accuracy, 1 percentage point higher sensitivity, and 7 percentage points lower mean unfairness (a 78% reduction) compared to the baseline, including reducing unfairness for Black and Hispanic non-white patients to 0. Thus, these estimators provided no appreciable loss in overall fit with only improvements in fairness. However, across most estimators, substantial unfairness remained for Asian patients, the smallest group we considered in our data sample, comprising only 2% of patients. The fair reductions constrained regression comparison estimators either had extremely low sensitivity or extremely low accuracy even when they improved unfairness, which is inappropriate for this application (tisdale_cost-effectiveness_2022).
@@
6 DISCUSSION
Although this framework is general, we do not empirically demonstrate how it performs for other fairness definitions, estimator types, or outcomes. Future areas of work include adapting our framework for such cases. We also found that for small groups of interest, our proposed method may have high variance and could be outperformed by alternative fair regression methods. In our application, we highlight that there is no universal standard for how race is measured across clinics in the American Family Cohort. We also lacked the sample size to study fairness for some racial groups, including American Indian and Alaska Native and Native Hawaiian and Pacific Islander patients. Furthermore, fair regression is just one aspect of ethical machine learning and should be considered alongside the full machine learning pipeline in health care (chen_ethical_2021).
Acknowledgements
Data for this project were accessed using the Stanford Center for Population Health Sciences Data Core. The PHS Data Core is supported by a National Institutes of Health National Center for Advancing Translational Science Clinical and Translational Science Award (UL1TR003142) and from Internal Stanford funding. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
We wish to acknowledge and thank the ABFM PRIME Registry participating clinicians and the American Board of Family Medicine, without whom the American Family Cohort would not be possible.
We also thank members of the Health Policy and Data Science Lab for their feedback on this research with special thanks to Malcolm Barrett for expert code review.
Funding
This project was supported by NIH grants R01LM013989 and 5T32HS026128-07, NSF grant DGE-2146755, and AHRQ grant 5T32HS026128-07.
Disclosures
The authors have no conflicts of interest to disclose. The project falls under Stanford IRB 68501, which approved human subjects involvement and waived consent.
Data availability statement
Due to ethical and legal restrictions on the use of patient health information, supporting data are not available.

### v2

@@ 4.1 Simulation Implementation
We aimed to explore the performance of various score functions constructed using each of the three group synthesis mechanisms in Section 3.2 with fairness score weights $\alpha\in\{0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,0.99\}$. Additionally, we studied different numbers of candidate penalty weight set draws $10,20,\dots,100]$ to empirically characterize changes in accuracy and fairness as the number of candidate sets increases. Lastly, we investigated robustness to functional form misspecification by omitting $X_{1}$ as an input variable. Comparison estimators were a baseline unpenalized regression and the fair reductions constrained regression method from Agarwal et al. (2018) as implemented in the fairlearn python package. The exponentiated gradient algorithm was implemented with allowed fairness constraint violations of $0.1$ and $0.01$, and the grid search algorithm was used with constraint weights of $0.25$, $0.5$, and $0.75$. Otherwise, default hyperparameters were used.
@@ 4.2 Simulation Results
Our novel penalized regression methods were evaluated across several key dimensions: behavior of unfairness penalty weights by fairness score weights, understanding penalty weights across various evaluation measures, effect of changes in multi-group unfairness synthesis function, and the tradeoff between fairness and accuracy. First, we characterized the behavior of unfairness penalty weights under the averaged score functions. The random search yielded different unfairness penalty weights for the three groups across the user-specified fairness score weights (Figure 1). Unfairness penalty weights increased, on average, as the fairness score weight $\alpha$ increased. Unfairness penalty weights also varied across groups and simulation settings.
Next, we examine the impact of the choice of multi-group synthesis function on accuracy and the fairness measures. All three methods for synthesis of unfairness measures across groups in the score function — population-weighted mean, group-weighted mean, and maximum — yielded similar results with overlapping interquartile ranges, especially for accuracy and population-weighted mean fairness (Figure 3). The methods only diverged in maximum unfairness, where the population-weighted mean method underperformed and the maximum method overperformed.
We compare the fairness and accuracy results of our penalized method to the fair reductions constrained regression method (Figure 4). This figure also shows the fairness-accuracy frontier, the set of results for which there is no better performance in both fairness and accuracy (Liang et al., 2023). The fairness-accuracy frontier for our penalized method was beyond that of the comparison fair reduction methods; the comparison method did not yield estimators that outperformed our method’s estimators in both mean fairness and accuracy. In simulation settings 1 and 2, the penalized methods were clustered around the fairness-accuracy frontier, while the comparison method had more variance. For simulation setting 3, which featured smaller groups, there was more variance across simulations for both methods, with more replications far from the fairness-accuracy frontier. Results for simulation performance studying variation in the number of candidate penalty weight sets and under estimator misspecification are available in Web Figures 1-2. Accuracy and unfairness had less variation as the number of candidate penalty weight sets increased and misspecified estimators generally had fairness-accuracy frontiers shifted downward.
5 CHRONIC KIDNEY DISEASE APPLICATION
5.1 Data
We applied our methods in the American Family Cohort dataset derived from the American Board of Family Medicine PRIME Registry (Phillips, 2017; Stanford Center for Population Health Sciences, 2024). These primary care practice electronic health record data from across the United States contain billing information (including diagnosis and procedure codes), drug information, and physiological patient measurements from laboratory tests and direct clinical observation. Using input variables measured in 2017-2018, we predicted the occurrence of ESRD in 2019. This outcome was defined through either an ESRD diagnosis code or serum creatinine corresponding to eGFR under 15. eGFR was calculated with the CKD-EPI 2021 equation, which does not adjust for race, to more accurately reflect the true physiological state of patients (Inker et al., 2021).
Predictors were selected based on prior literature (Segal et al., 2020; Inoguchi et al., 2022; Su et al., 2022; Wang et al., 2022). We had several physiologic predictors. The means of diastolic and systolic blood pressure as well as the maximum and mean values for hemoglobin A1C and billirubin were included as predictors, as was the minimum eGFR value. Missingness was accounted for through the inclusion of indicators of missingness as predictors. Indicators for hypertension emergencies, type 2 diabetes, glomerulonephritis, proteinuria, and pyelonephritis, constructed using diagnosis codes, were included as predictors as well. Diagnosis codes used for these measures are included in Web Table 1. Finally, binary documented sex, year of birth, race, and ethnicity were available in the American Family Cohort dataset. We included documented sex and age as predictors. Race and ethnicity were used to define the groups of interest in the construction of fairness penalties and metrics, but we intentionally did not include these variables as predictors. We studied fairness for three race and ethnicity groups: Black, Hispanic non-white, and Asian, with some patients identified as both Black and Hispanic non-white as well as Asian and Hispanic non-white. White patients were treated as the reference group. Two groups who face health disparities in CKD had small sample sizes (American Indian and Alaska Native as well as Native Hawaiian and Pacific Islander) in our dataset and were not included as groups for fairness optimization.
We restricted our sample to patients with diastolic and systolic blood pressure measurements in 2017-2018 and to practices with at least one ESRD patient in 2017-2019 for data quality. To target high-risk susceptible patients, the sample was restricted to patients at least 35 years old with at least two chronic kidney disease diagnosis codes, no documented ESRD prior to 2019, and a maximum systolic blood pressure of at least 140 mmHg, indicating stage 2 hypertension, in 2017-2018 (Hsu et al., 2005; Cusick et al., 2025). Finally, we required recorded values of either race or ethnicity. The resulting sample had 67397 patients, 2804 of whom developed ESRD in 2019. Detailed sample characteristics are described in Web Table 2.
5.2 Implementation
We used a decision threshold of 0.15 for estimated ESRD probability, informed by evidence from prior settings that the benefits of increasing the intensity of CKD treatment outweigh the costs and thus that false negatives are more harmful than false positives, but we recognize that other thresholds could also be valid (Tisdale et al., 2022). Penalty weight candidates were sampled from the distribution $\text{LogUnif}(-4,-1)$. We used three-fold cross validation to simultaneously identify the penalty weight sets that maximized each score function variation described in Section 4.1. Comparison estimators were the same as in our simulation with the same algorithms and hyperparameters described in Section 4.
5.3 Results
Under the 30 different score function formulations, 6 distinct sets of penalty weights that maximized the score functions and their corresponding penalized regression estimators were identified (Table 2). In most cases, more than one combination of synthesis method and fairness score weight yielded identical results. Overall, the penalized estimators had 0-4 percentage point decreases in accuracy with 1-7 percentage point decreases in unfairness. The penalized regression estimator with group-weighting and fairness score weight 0.1 had the same performance as the population-weighted estimator with fairness score weight 0.1 or 0.2. These three estimators had less than a percentage point difference in accuracy, 1 percentage point higher sensitivity, and 7 percentage points lower mean unfairness (a 78% reduction) compared to the baseline, including reducing unfairness for Black and Hispanic non-white patients to 0. Thus, these estimators provided no appreciable loss in overall fit with only improvements in fairness. However, across most estimators, substantial unfairness remained for Asian patients, the smallest group we considered in our data sample, comprising only 2% of patients. The fair reductions constrained regression comparison estimators either had extremely low sensitivity or extremely low accuracy even when they improved unfairness, which is inappropriate for this application (Tisdale et al., 2022).
6 DISCUSSION
Although this framework is general, we do not empirically demonstrate how it performs for other fairness definitions, estimator types, or outcomes. Future areas of work include adapting our framework for such cases. We also found that for small groups of interest, our proposed method may have high variance and could be outperformed by alternative fair regression methods. In our application, we highlight that there is no universal standard for how race is measured across clinics in the American Family Cohort. We also lacked the sample size to study fairness for some racial groups, including American Indian and Alaska Native and Native Hawaiian and Pacific Islander patients. Furthermore, fair regression is just one aspect of ethical machine learning and should be considered alongside the full machine learning pipeline in health care (Chen et al., 2021).
Acknowledgements
Data for this project were accessed using the Stanford Center for Population Health Sciences Data Core. The PHS Data Core is supported by a National Institutes of Health National Center for Advancing Translational Science Clinical and Translational Science Award (UL1TR003142) and from Internal Stanford funding. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH. We wish to acknowledge and thank the ABFM PRIME Registry participating clinicians and the American Board of Family Medicine, without whom the American Family Cohort would not be possible. We also thank members of the Health Policy and Data Science Lab for their feedback on this research with special thanks to Malcolm Barrett for expert code review.
This project was supported by NIH grant R01LM013989, NSF grant DGE-2146755, and AHRQ grant 5T32HS026128-07. The authors have no conflicts of interest to disclose. The project falls under a Stanford IRB, which approved human subjects involvement and waived consent.
Data availability statement
The data underlying this article cannot be shared publicly due to data use agreement restrictions.
@@ Acknowledgements
Supplementary Material
Web Appendices, Tables, and Figures, and code referenced in Sections 3, 3.1, 4, 4.2, and 5.1 are available below. Code is available in a public GitHub repository at the link https://github.com/StanfordHPDS/penalized_multiple_groups.
Web Appendix A: Continuous Outcomes
Here we seek to estimate continuous outcome $Y$. The baseline continuous outcome regression generates estimates $\hat{Y}$ as a function of $m$ input variables $X$, indexed by $q$, by identifying the estimator that minimizes the mean squared error (MSE) loss. The unfairness metric is mean residual disparity, defined as the difference in mean residuals between the reference group and the group of interest:
$U_{MR}=\frac{1}{n_{r}}\sum_{k\in r}(\hat{Y}_{k}-Y_{k})-\frac{1}{n_{g}}\sum_{j\in g}(\hat{Y}_{j}-Y_{j}).$} (3)
Penalized regression with continuous outcomes will add penalties for unfairness as defined in Equation 3, where each group has its own penalty term and penalty weight.
Example: Lasso
We first use the example of lasso, which involves adding a penalty on regression coefficients $\beta$ with penalty weight $\gamma$ to the mean squared error loss and estimating $\hat{Y}$ as a linear combination of regression coefficients and predictors:
$L=\sum_{i=1}^{n}\left[\sum_{q=1}^{m}\beta_{q}X_{iq}-Y_{i}\right]^{2}+\gamma\sum_{q=1}^{m}|\beta_{q}|.$
The loss function with fairness penalties is given by:
$L=\sum_{i=1}^{n}\left[\sum_{q=1}^{m}\beta_{q}X_{iq}-Y_{i}\right]^{2}+\gamma\sum_{q=1}^{m}|\beta_{q}|+\sum_{g\in G}\lambda_{g}\left[\frac{1}{n_{r}}\sum_{k\in r}(\hat{Y}_{k}-Y_{k})-\frac{1}{n_{g}}\sum_{j\in g}(\hat{Y}_{j}-Y_{j})\right].$
This loss function minimization problem can be solved using the L-BFGS algorithm, a type of quasi-Newtonian optimizer (Liu and Nocedal, 1989), available in the PyTorch package.
Example: Neural Net
A penalized loss function for a simple neural network with two hidden layers could be defined as:
$L=\sum_{i=1}^{n}\left[\sum_{q=1}^{m}\beta_{q}X_{iq}-Y_{i}\right]^{2}+\sum_{g\in G}\lambda_{g}\left[\frac{1}{n_{r}}\sum_{k\in r}(\hat{Y}_{k}-Y_{k})-\frac{1}{n_{g}}\sum_{j\in g}(\hat{Y}_{j}-Y_{j})\right].$} (4)
The loss function minimization problem here can be solved using the Adam algorithm, a type of gradient descent-based optimizer (Loshchilov and Hutter, 2019), also available in the PyTorch package.
Web Appendix B: Weighted Classification Derivation
We begin with the loss function from section 3.1 of the manuscript.
$L=-\sum_{i=1}^{n}\{Y_{i}\ln(\hat{P}_{i})+(1-Y_{i})\ln(1-\hat{P}_{i})\}+\sum_{g\in G}\lambda_{g}n_{g}\left(\frac{\sum_{k\in r}Y_{k}\hat{P}_{k}}{\sum_{k\in r}Y_{k}}-\frac{\sum_{j\in g}Y_{j}\hat{P}_{j}}{\sum_{j\in g}Y_{j}}\right).$
We could calculate the gradient and use a gradient descent-based algorithm to find the solution. With $\nabla\hat{P}=\hat{P}(1-\hat{P})X$, the gradient is
However, there is a more efficient method available based on the reductions approach of Agarwal et al. (2018). In this approach, we use cost minimization loss, which punishes misalignment between $\hat{P}_{i}$ and $Y_{i}$ with a cost and minimizes total costs (Zadrozny et al., 2003). The cost-sensitive loss function with the standard erroneous classification costs of 1 for all observations, which has the same solution as the standard binary cross entropy loss, can be added to the true positive rate disparity unfairness penalty terms to get the cost minimization penalized loss function:
$L=\sum_{i=1}^{n}\left\{\hat{P}_{i}(1-Y_{i})+(1-\hat{P}_{i})Y_{i}\right\}+\sum_{g\in G}\lambda_{g}n_{g}\left(\frac{\sum_{k\in r}Y_{k}\hat{P}_{k}}{\sum_{k\in r}Y_{k}}-\frac{\sum_{j\in g}Y_{j}\hat{P}_{j}}{\sum_{j\in g}Y_{j}}\right).$
$L=\sum_{i=1}^{n}\left[\hat{P}_{i}\left\{1-Y_{i}+Y_{i}\sum_{g\in G}\lambda_{g}n_{g}\left(\frac{\mathbb{I}_{ri}}{\sum_{k\in r}Y_{k}}-\frac{\mathbb{I}_{gi}}{\sum_{j\in g}Y_{j}}\right)\right\}+(1-\hat{P}_{i})Y_{i}\right].$
With this formulation, we extend a theorem from Zadrozny et al. (2003) to rewrite the cost-sensitive classification problem as a weighted non-cost-sensitive classification problem. In this formulation, the weights are
$W_{i}=|C_{i}^{0}-C_{i}^{1}|=\left|Y_{i}\left\{2-\sum_{g\in G}\lambda_{g}n_{g}\left(\frac{\mathbb{I}_{ri}}{\sum_{k\in r}Y_{k}}-\frac{\mathbb{I}_{gi}}{\sum_{j\in g}Y_{j}}\right)\right\}-1\right|,$
Web Appendix C: Simulation Study
Let $g_{A}$, $g_{B}$, and $g_{C}$ be the three different groups whose membership was reflected by indicators $\mathbb{I}_{g_{A}}$, $\mathbb{I}_{g_{B}}$, and $\mathbb{I}_{g_{C}}$, respectively. We defined the reference group to be the set of individuals who were not in any of the three groups. The data were generated as follows with $n=1,000,000$:
$X_{1},X_{2}\sim\mathcal{N}(30,15)$
$X_{3},X_{4}\sim Pois(15)$
$X_{5},\dots,X_{9}\sim Bern(p),p=[0.2,0.4,0.6,0.8,0.95]$
$\mathbb{I}_{g_{A}}\sim Bern(\min[\max\{c_{0}(X_{9}+0.1)+c_{3}X_{6}+\mathcal{N}(0,0.02),0\},1])$
$\mathbb{I}_{g_{B}}\sim Bern(\min[\max\{c_{1}(X_{7}+0.05)+c_{4}X_{6}+\mathcal{N}(0,0.02),0\},1])$
$\mathbb{I}_{g_{C}}\sim Bern(\min[\max\{c_{2}(X_{8}+0.4)+c_{5}X_{6}+\mathcal{N}(0,0.02),0\},1])$
$Y$ $\sim Bern\Bigg(\Phi^{-1}\bigg[\frac{X_{1}}{30}-\frac{X_{2}}{15}+\frac{3X_{3}}{50}-\frac{X_{4}}{25}+\frac{1}{100}\Big\{2^{X_{5}+X_{6}}+6X_{7}$
[…]
Web Table 1: Identification Strategy for ABFM Variables. For most variables, we queried for the enumerated OMOP concepts in the enumerated fields. For other variables, we queried the OMOP concept lookup table for the OMOP ABFM data with the given text.
Diagnosis Search Method Query
ESRD Text Query “End stage renal disease” or “End stage renal failure”
eGFR Preexisting See repository
HbA1c OMOP Concepts: Measurement 4184637, 36304734
Bilirubin OMOP Concepts: Measurement 4118986, 3024148, 4216632, 3024128
Diastolic BP OMOP Concepts: Measurement 4298393, 3012888
Systolic BP OMOP Concepts: Measurement 4152194, 3004249
CKD OMOP Concepts: Condition 44784621, 443611, 443601, 44782690, 43021852, 45768812, 44782728, 45763855, 45763854, 43531578, 45768812, 443612, 44782429, 43531653, 443614, 43021852, 443597
Diabetes (Type 2) OMOP Concepts: Condition / ICD10 44809548, 37208172 / E11
Kidney Stone OMOP Concepts: Condition 201620, 45548655, 44822046
Obstruction OMOP Concepts: Condition 45612141, 45592125, 45567838, 36713449, 4220439, 433813
Hypertension Emergency OMOP Concepts: Condition 43020424, 45768449, 37200492, 37200493, 37200494
Proteinuria Text Query “Proteinuria”
Glomerulonephritis Text Query “Glomerulonephritis”
Pyelonephritis Text Query “Pyelonephritis”
Female OMOP Concepts: Gender 8532
White OMOP Concepts: Race 8527, 38003614
Black OMOP Concepts: Race 8516, 38003597, 38003598, 38003599, 38003600
Asian OMOP Concepts: Race 8515, 38003574-38003596
Hispanic OMOP Concepts: Ethnicity 38003563
Web Table 2: Cohort Patient Characteristics in 2017-2018 by ESRD Status in 2019.
Missing Overall No ESRD ESRD
n 0 67397 64593 2804
Measurements
Minimum eGFR, mean (SD) 36899 48.3 (18.7) 49.5 (18.0) 22.3 (16.1)
Mean Diastolic BP, mean (SD) 0 75.3 (7.9) 75.4 (7.9) 74.2 (8.8)
Mean Systolic BP, mean (SD) 0 136.6 (11.9) 136.5 (11.7) 140.0 (14.1)
Mean HbA1c, mean (SD) 44250 6.8 (1.4) 6.8 (1.4) 6.9 (1.6)
Mean Bilirubin, mean (SD) 34493 0.6 (0.6) 0.6 (0.6) 0.5 (0.3)
Diagnoses
Diabetes (Type 2), n (%) 0 47447 (70.4) 45780 (70.9) 1667 (59.5)
Kidney Stone, n (%) 0 32734 (48.6) 30947 (47.9) 1787 (63.7)
Proteinuria, n (%) 0 1255 (1.9) 1223 (1.9) 32 (1.1)
Pyelonephritis, n (%) 0 494 (0.7) 451 (0.7) 43 (1.5)
Demographics
Age, mean (SD) 0 75.0 (11.3) 75.1 (11.3) 72.0 (12.3)
Female, n (%) 0 35655 (52.9) 34283 (53.1) 1372 (48.9)
White, n (%) 0 53402 (79.2) 51519 (79.8) 1883 (67.2)
Black, n (%) 0 8094 (12.0) 7558 (11.7) 536 (19.1)
Asian, n (%) 0 1427 (2.1) 1325 (2.1) 102 (3.6)
Hispanic non-white, n (%) 0 2678 (4.0) 2476 (3.8) 202 (7.2)
(a) Accuracy
(b) Population-Weighted Unfairness
Web Figure 1: Fair Regression Estimator Accuracy and Unfairness by Number of Candidate Unfairness Penalty Weight Draws Across Different Simulation Settings. Confidence intervals represent the empirical 25th and 75th percentiles of replications.
(a) Simulation Setting 1
(b) Simulation Setting 2
(c) Simulation Setting 3
Web Figure 2: Fair Regression Estimator Accuracy and Fairness for Correctly Specified and Misspecified Fair Regression Estimators by Simulation Setting. Each plotted point corresponds to the result of a single simulation experiment under one of the score function variations.